OpenStudy (anonymous):

Need help!!! Find a quadratic function whose x-intercepts are -4 and 2 and whose range is [-18, 00).

OpenStudy (anonymous):

So we know that it has x-intercepts of -4 and 2 and that it's vertex is in between those two at a y-value of -18! The average of the two is: \[x_{av}=\frac{x_1+x_2}{2}=\frac{-4+2}{2}=\frac{-2}{2}=-1\] So then, we know that the vertex is located at the point of \(P(-1,-18)\) In general, for a parabola with the vertex located at the point \(V(h,k)\), the formula is: \[f(x)=a(x-h)^2+k\] So we know: \(h=-1\) \(k=-18\) And we can simplify the equation: \[\eqalign{ &f(x)=a(x-(-1))^2+(-18) \\ &f(x)=a(x+1)^2-18 }\] Now we need to find \(a\). Following so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but how do we find a?

OpenStudy (jdoe0001):